OpenStudy (anonymous):

Solve. x^4+8x^3+21x^2+18x is less than or equal to 0

OpenStudy (anonymous):

\[x^4+8x^3+21x^2+18x \le 0\]

OpenStudy (anonymous):

Notice that \[ x^4+8 x^3+21 x^2+18 x=x (x+2) (x+3)^2 \] It is enough to see when \[ x (x+2) <0 \]

OpenStudy (anonymous):

This happens if -2 < x < 0

OpenStudy (anonymous):

how did you factor it like that?

OpenStudy (anonymous):

and how come (x+3)^2 is not less than 0?